# Learning Robust and Correct Controllers Guided by Feasibility-Aware Signal Temporal Logic via BarrierNet

### v2

@@ V-A2 Neural Network Controller with Memory
$1-\sqrt[4]{(\frac{x-o_{x,i}}{a_{i}})^{4}+(\frac{y-o_{y,i}}{b_{i}})^{4}}\geq 0,\ j=3,4,5.$ (47)
Here, $Reg_{j}$ belongs to Category II and $Obs_{j}$ belongs to Category I for $j=1,2,3,4,5$. In plain terms, the STL formula $\varphi$ requires the robot to reach $Reg_{1}$ within the time interval $[0,5]$ and $Reg_{2}$ within $[5,10]$, while avoiding obstacles $Obs_{3}$, $Obs_{4}$ and $Obs_{5}$ at all times. The time horizon of $\varphi$ is $10$. All five predicates have a relative degree of $2$ with respect to system (41). In this example, we set $\mathbf{Q}(\cdot,\bm{\theta}_{q})$ to be an identity matrix, so for all $t>0$ the output of the previous layers reduces to $\mathbf{F}(\cdot,\bm{\theta}_{f})$, which serves as a reference control. Because the task requires the robot to visit $Reg_{1}$ and then return to $Reg_{2}$ (within the $Init$ region), the motion involves a back-and-forth behavior. Therefore, we evaluate both memory-based and memoryless structures for generating $\mathbf{F}$. For the memoryless structure, RefNet is implemented as a feedforward neural network with three fully connected layers. For the memory-based structure, RefNet is implemented as a 2-layer recurrent neural network (LSTM). In both cases, InitNet and BarrierNet are implemented as feedforward neural networks with three fully connected layers each. For the robustness measure, we adopt the exponential robustness in (29) and (30).
@@
Figure 7: Inputs over time for 4 trajectories from sampled initial conditions under FeasiBN-VarP (solid curves: controller with memory. Dashed curves: controller without memory).
@@ V-B Case Study II: Nonlinear Dynamical System — Unicycle
where $\mathbf{x}=[x\ y\ \theta\ v]^{\top}$ and $\mathbf{u}=[u_{1}\ u_{2}]^{\top}$, with $x\ y]^{\top}$ denoting the 2D position, $\theta$ the heading angle, $v$ the linear speed, $u_{1}$ the angular velocity, and $u_{2}^{\top}$ the acceleration of the robot. During the initialization of the neural network controller, the control inputs are bounded. However, in the subsequent optimization stage, the inputs are treated as unconstrained, while the STL specification still accounts for the actual input limits in (34), and the cost function in (40) employs an $\ell_{2}$ penalty with a weight of $0.003$ to discourage large accelerations.
@@
$1-\sqrt[4]{(\frac{x-o_{x,i}}{a_{i}})^{4}+(\frac{y-o_{y,i}}{b_{i}})^{4}}\geq 0,\ j=3,4,5.$ (51)
Here, $Reg_{j}$ belongs to Category II and $Obs_{j}$ belongs to Category I for $j=1,2,3,4,5$. In plain terms, the STL formula $\varphi$ requires the robot to reach $Reg_{1}$ within the time interval $[0,3]$ and $Reg_{2}$ within $[3,6]$, while avoiding obstacles $Obs_{3}$, $Obs_{4}$ and $Obs_{5}$ at all times. The time horizon of $\varphi$ is $6$. All five predicates have a relative degree of $2$ with respect to system (41). In this example, we set $\mathbf{Q}(\mathbf{x}(t),\bm{\theta}_{q})$ to be an identity matrix, so for all $t>0$ the output of the previous layers reduces to $\mathbf{F}(\mathbf{x}(t),\bm{\theta}_{f})$, which serves as a memoryless reference control. We directly adopt the FeasiBN-VarP configuration described in Sec. V-A1. RefNet and InitNet are implemented as feedforward neural networks with three fully connected layers each. For the robustness measure, we adopt the exponential robustness in (29) and (30).
@@ VI Conclusion and Future Work
Shuo Liu (Student Member, IEEE) received his B.Eng. degree in Mechanical Engineering from Chongqing University, Chongqing, China, in 2018, and his M.Sc. degree in Mechanical Engineering from Columbia University, New York, NY, USA, in 2020. He is currently a Ph.D. candidate in Mechanical Engineering at Boston University, Boston, USA and his research interests include optimization, nonlinear control, deep learning and robotics.
Wenliang Liu (Member, IEEE) received the B.Sc. degree in instrumentation science from Beihang University, Beijing, China, in 2019, and the M.Sc. and Ph.D. degrees in mechanical engineering from Boston University, Boston, MA, USA, in 2023 and 2024. His research interests include control theory, formal methods, and machine learning, with robotics applications.
Wei Xiao (Member, IEEE) received the B.Sc. degree in mechanical engineering and automation from the University of Science and Technology Beijing, Beijing, China, the M.Sc. degree in robotics from the Chinese Academy of Sciences (Institute of Automation), Beijing, China, and the Ph.D. degree in systems engineering from Boston University, Boston, MA, USA, in 2013, 2016, and 2021, respectively.
He is currently an Assistant Professor at Worcester Polytechnic Institute (WPI), Worcester, MA, USA. His current research interests include control theory and machine learning, with particular emphasis on robotics and traffic control.
Calin Belta (Fellow, IEEE) received the B.Sc. and M.Sc. degrees from the Technical University of Iasi, Iasi, Romania, in 1995 and 1997, respectively, and the M.Sc. and Ph.D. degrees from the University of Pennsylvania, Philadelphia, PA, USA, in 2001 and 2003.
He is currently the Brendan Iribe Endowed Professor of Electrical and Computer Engineering and Computer Science at the University of Maryland, College Park, MD, USA. His research focuses on control theory and formal methods, with particular emphasis on hybrid and cyber-physical systems, synthesis and verification, and applications in robotics and biology.
Dr. Belta was a recipient of the Air Force Office of Scientific Research Young Investigator Award and the National Science Foundation CAREER Award.

### v3

@@ V-A2 Neural Network Controller with Memory
where $\mathbf{u}_{\min}=[-10\ -10]^{\top},$ and $\mathbf{u}_{\max}=[10\ 10]^{\top}.$ $Obs_{j}$ is a superellipse defined by (43) where $j=3,4,5$. Here, $Reg_{j}$ belongs to Category II and $Obs_{j}$ belongs to Category I for $j=1,2,3,4,5$. In plain terms, the STL formula $\varphi$ requires the robot to reach $Reg_{1}$ within the time interval $[0,5]$ and $Reg_{2}$ within $[5,10]$, while avoiding obstacles $Obs_{3}$, $Obs_{4}$ and $Obs_{5}$ at all times. The time horizon of $\varphi$ is $10$. All five predicates have a relative degree of $2$ with respect to system (40). In this example, we set $\mathbf{Q}(\cdot,\bm{\theta}_{q})$ to be an identity matrix, so for all $t>0$ the output of the previous layers reduces to $\mathbf{F}(\cdot,\bm{\theta}_{f})$, which serves as a reference control. Because the task requires the robot to visit $Reg_{1}$ and then return to $Reg_{2}$ (within the $Init$ region), the motion involves a back-and-forth behavior. Therefore, we evaluate both memory-based and memoryless structures for generating $\mathbf{F}$. For the memoryless structure, RefNet is implemented as a feedforward neural network with three fully connected layers. For the memory-based structure, RefNet is implemented as a 2-layer recurrent neural network (LSTM). In both cases, InitNet and BarrierNet are implemented as feedforward neural networks with three fully connected layers each. For the robustness measure, we adopt the exponential robustness in (29) and (30).
@@
Figure 7: Inputs over time for 4 trajectories from sampled initial conditions under FeasiBN-VarP (solid curves: controller with memory; dashed curves: controller without memory).
@@ V-B Case Study II: Nonlinear Dynamical System — Unicycle
where $\mathbf{x}=[x\ y\ \theta\ v]^{\top}$ and $\mathbf{u}=[u_{1}\ u_{2}]^{\top}$, with $x\ y]^{\top}$ denoting the 2D position, $\theta$ the heading angle, $v$ the linear speed, $u_{1}$ the angular velocity, and $u_{2}^{\top}$ the acceleration of the robot. The control input treatment and cost function settings follow those in Sec. V-A.
@@
where $\mathbf{u}_{\min}=[-10\ -10]^{\top},$ and $\mathbf{u}_{\max}=[10\ 10]^{\top}$. $Obs_{j}$ is a superellipse defined by (43) where $j=3,4,5$. Here, $Reg_{j}$ belongs to Category II and $Obs_{j}$ belongs to Category I for $j=1,2,3,4,5$. In plain terms, the STL formula $\varphi$ requires the robot to reach $Reg_{1}$ within the time interval $[0,3]$ and $Reg_{2}$ within $[3,6]$, while avoiding obstacles $Obs_{3}$, $Obs_{4}$ and $Obs_{5}$ at all times. The time horizon of $\varphi$ is $6$. All five predicates have a relative degree of $2$ with respect to system (40). In this example, we set $\mathbf{Q}(\mathbf{x}(t),\bm{\theta}_{q})$ to be an identity matrix, so for all $t>0$ the output of the previous layers reduces to $\mathbf{F}(\mathbf{x}(t),\bm{\theta}_{f})$, which serves as a memoryless reference control. We directly adopt the FeasiBN-VarP configuration described in Sec. V-A1. RefNet and InitNet are implemented as feedforward neural networks with three fully connected layers each. For the robustness measure, we adopt the exponential robustness in (29) and (30).